# Character Complexity: A Novel Measure for Quantum Circuit Analysis

### v2

@@ 1 Introduction
Key Findings
  1. 1.
    Novel Complexity Measure: I introduce “Character Complexity” as a new measure for quantum circuit analysis, based on group-theoretic principles.
  2. 2.
    Theoretical Bounds: Character Complexity ($C(U)$) is bounded by $0\leq C(U)\leq\min(n,\operatorname{rank}(U))$, where $n$ is the number of qubits.
  3. 3.
    Composition Properties: Character Complexity is multiplicative under tensor products and submultiplicative under circuit composition.
  4. 4.
    Classical Simulation Threshold: Quantum circuits with $C(U)\leq c\log(n)$ can be efficiently simulated classically under certain conditions, providing a new criterion for classical-quantum boundaries.
  5. 5.
    Geometric Interpretation: I provide a novel visualization of Character Complexity using a modified Bloch sphere representation.
  6. 6.
    Scaling Behavior: Empirical results suggest Character Complexity scales as $O(n\log m)$ with $n$ qubits and $m$ gates, offering insights into quantum circuit structure.
  7. 7.
    Practical Implications: Character Complexity provides new perspectives on quantum algorithm design and optimization, potentially leading to more efficient quantum circuits.
@@ 3 Character Complexity: Definition and Properties
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$ (5)
@@
Let $U$ be a unitary operation and $G$ be a finite group. The character complexity $C(U)$ is bounded as follows:
$0\leq C(U)\leq\sqrt{\frac{k}{|G|}}\leq 1$
where $k$ is the number of irreducible representations of $G$.
@@
We prove each part of the inequality separately:
1. Lower bound: $0\leq C(U)$
This follows directly from the definition of $C(U)$:
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
Since $|\chi_{i}(U)|^{2}\geq 0$ and $d_{i}>0$ for all $i$, each term in the sum is non-negative, and thus $C(U)\geq 0$.
2. Upper bound: $C(U)\leq\sqrt{\frac{k}{|G|}}\leq 1$
We start by applying the Cauchy-Schwarz inequality:
$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}$ $\leq\sqrt{\sum_{i}|\chi_{i}(U)|^{2}}\sqrt{\sum_{i}\frac{1}{d_{i}}}$
We use the character orthogonality relation [11]:
$\sum_{i}|\chi_{i}(U)|^{2}=|G|\cdot|\{g\in G:gU=Ug\}|$
Next, we apply the Cauchy-Schwarz inequality to the dimension sum formula $\sum_{i}d_{i}^{2}=|G|$ [10]:
$\sum_{i}\frac{1}{d_{i}}\leq\sqrt{\sum_{i}1}\sqrt{\sum_{i}\frac{1}{d_{i}^{2}}}=\sqrt{k}\sqrt{\frac{1}{|G|}}$
Combining these results:
$C(U)$ $=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
$\leq\frac{1}{|G|}\sqrt{|G|\cdot|\{g\in G:gU=Ug\}|}\sqrt{k}\sqrt{\frac{1}{|G|}}$
=|$=\sqrt{\frac{|\{g\in G:gU=Ug\}|}{|G|}}\sqrt{\frac{k}{|G|}}$
Since |$|\{g\in G:gU=Ug\}|$ is the order of a subgroup of $G$, we have |$|\{g\in G:gU=Ug\}|\leq|G|$. Therefore:
$C(U)\leq\sqrt{\frac{k}{|G|}}$
Finally, since $k\leq|G|$ for any finite group (as the number of irreducible representations cannot exceed the order of the group), we have:
$\sqrt{\frac{k}{|G|}}\leq\sqrt{\frac{|G|}{|G|}}=1$
Thus, the theorem is proved. ∎
Remark 1.
The character complexity $C(U)$ measures how ”spread out” the action of $U$ is across different irreducible representations of the group $G$. A low value of $C(U)$ suggests that $U$’s action is concentrated in a few representations, indicating a simpler or more structured operation. Conversely, a value of $C(U)$ approaching $\sqrt{\frac{k}{|G|}}$ indicates that $U$’s action is more evenly distributed across many representations, suggesting a more complex or less structured operation.
Values will generally fall between 0 and $\sqrt{\frac{k}{|G|}}$, with values close to $\sqrt{\frac{k}{|G|}}$ indicating operations that maximally spread their action across all irreducible representations, and values close to 0 suggesting operations that act more trivially. This measure provides a way to quantify the complexity of operations that goes beyond simple counts, offering insights into the underlying structure of algorithms by emphasizing the relative distribution of an operation’s action across the irreducible representations.
@@ 1. Lower bound: $0\leq C(U)$
Theorem 3 (Submultiplicativity under Composition).
For quantum circuits $U_{1}$ and $U_{2}$ acting on the same set of qubits:
$C(U_{1}U_{2})\leq C(U_{1})\cdot C(U_{2})$
@@
Let $G$ be the finite group generated by the gates in both $U_{1}$ and $U_{2}$.
1) Recall the definition of character complexity:
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right)$
2) For the composition $U_{1}U_{2}$, we can use the Cauchy-Schwarz inequality:
$|\chi_{i}(U_{1}U_{2})|^{2}\leq|\chi_{i}(U_{1})|^{2}\cdot|\chi_{i}(U_{2})|^{2}$
3) Applying this to the character complexity:
$C(U_{1}U_{2})$ $=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1}U_{2})|^{2}}{d_{i}}\right)$
$\leq\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1})|^{2}\cdot|\chi_{i}(U_{2})|^{2}}{d_{i}}\right)$
4) Now, we can apply the Cauchy-Schwarz inequality again to the sum:
$\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1})|^{2}\cdot|\chi_{i}(U_{2})|^{2}}{d_{i}}\right)$ $\leq\sqrt{\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{1})|^{2}}{d_{i}}\right)}\cdot\sqrt{\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U_{2})|^{2}}{d_{i}}\right)}$
$=\sqrt{C(U_{1})}\cdot\sqrt{C(U_{2})}$
5) Finally, since $C(U)\geq 1$ for all unitary operations $U$, we have:
$\sqrt{C(U_{1})}\cdot\sqrt{C(U_{2})}\leq C(U_{1})\cdot C(U_{2})$
Therefore, we have shown that:
$C(U_{1}U_{2})\leq C(U_{1})\cdot C(U_{2})$
Thus, the theorem is proved. ∎
@@ 7 Discussion
One of the key insights gained from character complexity analysis is the existence of a tight upper bound on the complexity of a quantum circuit based on the number of qubits and the rank of the circuit’s matrix representation. This bound provides a fundamental limit on the complexity of quantum circuits and could guide the design of optimal quantum algorithms.

### v3

@@ 3 Character Complexity: Definition and Properties
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right),$ (5)
@@
Let $U$ be a unitary operation acting on $n$ qubits, and let $G$ be a finite group. The character complexity $C(U)$ is bounded as follows:
$0\leq C(U)\leq 1$
@@
1. Non-negativity: $C(U)\geq 0$
By definition,
$C(U)=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\right).$
Since $|\chi_{i}(U)|^{2}\geq 0$ and $d_{i}>0$ for all $i$, each term $\frac{|\chi_{i}(U)|^{2}}{d_{i}}\geq 0$. Therefore, the sum and hence $C(U)$ are non-negative:
$C(U)\geq 0.$
2. Upper Bound: $C(U)\leq 1$
First, recall that for any unitary representation, the absolute value of the character is bounded by the dimension of the representation:
$|\chi_{i}(U)|\leq d_{i}.$ (6)
This is because the character $\chi_{i}(U)$ is the trace of the unitary matrix representing $U$ in the $i$-th irreducible representation, and the absolute value of the trace of a unitary matrix is bounded by its dimension.
Squaring both sides, we get:
$|\chi_{i}(U)|^{2}\leq d_{i}^{2}.$ (7)
Therefore,
$\frac{|\chi_{i}(U)|^{2}}{d_{i}}\leq\frac{d_{i}^{2}}{d_{i}}=d_{i}.$ (8)
Now, summing over all $i$:
$\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}}\leq\sum_{i}d_{i}.$ (9)
Note that the sum of the dimensions of all irreducible representations satisfies:
$\sum_{i}d_{i}=|G_{\text{conj}}|,$ (10)
where $|G_{\text{conj}}|$ is the number of conjugacy classes in $G$. Since $|G_{\text{conj}}|\leq|G|$, we have:
$\sum_{i}d_{i}\leq|G|.$ (11)
However, to ensure an exact upper bound, consider that the maximum value of $C(U)$ occurs when $U$ is the identity element $e$ of $G$. In this case, the characters satisfy:
$\chi_{i}(e)=d_{i}.$ (12)
Therefore,
$C(e)$ $=\frac{1}{|G|}\left(\sum_{i}\frac{|\chi_{i}(e)|^{2}}{d_{i}}\right)$
$=\frac{1}{|G|}\left(\sum_{i}\frac{d_{i}^{2}}{d_{i}}\right)$
$=\frac{1}{|G|}\left(\sum_{i}d_{i}\right).$
Since $C(e)$ is maximized at the identity element, and $\sum_{i}d_{i}\leq|G|$, it follows that:
$C(U)\leq C(e)\leq 1.$ (13)
Therefore,
$0\leq C(U)\leq 1$ (14)
@@ 2. Upper Bound: $C(U)\leq 1$
Theorem 3 (General Bound on Character Complexity under Composition).
Let $G$ be a finite group, and let $U_{1},U_{2}\in G$. Then the character complexity of their composition satisfies:
$C(U_{1}U_{2})\leq 1.$
@@
For any $U\in G$, the character complexity is defined as:
$C(U)=\frac{1}{|G|}\sum_{i}\frac{|\chi_{i}(U)|^{2}}{d_{i}},$
where $\chi_{i}(U)$ is the character of $U$ in the $i$-th irreducible representation, and $d_{i}$ is its dimension.
Since $|\chi_{i}(U)|\leq d_{i}$ for all $i$ and $U\in G$, we have:
$\frac{|\chi_{i}(U)|^{2}}{d_{i}}\leq d_{i}.$
Therefore, for the composition $U_{1}U_{2}$:
$C(U_{1}U_{2})=\frac{1}{|G|}\sum_{i}\frac{|\chi_{i}(U_{1}U_{2})|^{2}}{d_{i}}\leq\frac{1}{|G|}\sum_{i}d_{i}=1,$
since $\sum_{i}d_{i}=|G|$. ∎
Lemma 1 (Multiplicativity of Character Complexity for Abelian Groups).
Let $G$ be a finite abelian group, and let $U_{1},U_{2}\in G$. Then the character complexity is multiplicative under composition:
$C(U_{1}U_{2})=C(U_{1})\cdot C(U_{2}).$
Proof.
In finite abelian groups, all irreducible representations are one-dimensional ($d_{i}=1$), and the characters satisfy:
$\chi_{i}(U_{1}U_{2})=\chi_{i}(U_{1})\chi_{i}(U_{2}).$
Therefore, the character complexity simplifies to:
$C(U)=\frac{1}{|G|}\sum_{i}|\chi_{i}(U)|^{2}.$
Computing the character complexity of the composition:
$C(U_{1}U_{2})$ $=\frac{1}{|G|}\sum_{i}|\chi_{i}(U_{1}U_{2})|^{2}$
$=\frac{1}{|G|}\sum_{i}|\chi_{i}(U_{1})\chi_{i}(U_{2})|^{2}$
$=\frac{1}{|G|}\sum_{i}|\chi_{i}(U_{1})|^{2}|\chi_{i}(U_{2})|^{2}$
$=\left(\frac{1}{|G|}\sum_{i}|\chi_{i}(U_{1})|^{2}\right)\left(\frac{1}{|G|}\sum_{i}|\chi_{i}(U_{2})|^{2}\right)$
$=C(U_{1})\cdot C(U_{2}).$
@@ 7 Discussion
One of the key insights gained from character complexity analysis is the existence of a tight upper bound on the complexity of a quantum circuit. This bound provides a fundamental limit on the complexity of quantum circuits and could guide the design of optimal quantum algorithms.